# Swin fMRI Transformer Predicts Early Neurodevelopmental Outcomes from Neonatal fMRI

### v1

@@ Results
For each dimensionality, the model was tuned using a predefined hyperparameter search space, including learning rate, patch size, and data augmentation strategies (see Section 3.3). With ICA reducing the input data to a fraction of its original dimensionality, careful tuning was critical to ensure effective convergence. Given the data demands of transformer architectures and the reduced training data size, we increased the maximum training epochs to 100. To mitigate overfitting, we applied data augmentation techniques such as affine and intensity transformations to enhance sample diversity and improve generalizability. Due to the specific sequence lengths required for IC-based training (42 and 100), transfer learning was not utilized in these experiments.
We addressed potential data leakage by ensuring that the 100 healthy subjects used for Group ICA were always included in the training set, avoiding overlap with validation or test sets. This step minimized bias and maintained the reliability of our results. Experimental outcomes are summarized in Tables 5 and 6, presenting the average test performance and standard deviations of the fine-tuned models.
@@ Discussion
This study demonstrates that integrating Group ICA-based dimensionality reduction with the SwiFT Transformer architecture significantly enhances the prediction of neurodevelopmental outcomes from neonatal fMRI data. By extracting biologically meaningful features through ICA and employing multi-label learning, our approach balances dimensionality reduction with the preservation of critical neural information. The SwiFT Transformer uniquely learns complex spatiotemporal patterns in early brain development directly from 4D fMRI data, leveraging its attention-based architecture to model both local and global dependencies. These findings highlight the potential of combining neuroscience-driven feature engineering with advanced machine learning to improve both interpretability and predictive accuracy, providing a robust framework for neurodevelopmental research and clinical applications.
Our results indicate that the choice of ICA dimensionality (25 vs. 100 components) does not significantly affect the model’s predictive performance in regression or classification tasks. This invariance suggests that ICA robustly captures the essential features of neural activity, regardless of the specific number of components used. Both configurations effectively isolate key brain networks while filtering out noise, ensuring that critical information relevant to neurodevelopmental outcomes is retained. This finding reinforces the adaptability of ICA-based dimensionality reduction as a reliable preprocessing step for fMRI analysis.
As shown in Table 5, single-label learning with ICA-extracted features slightly outperforms predictions using raw fMRI data, particularly for language skill classification and motor skill regression. These results demonstrate that dimensionality reduction through ICA retains critical neural information while effectively filtering out noise. By isolating independent components that likely represent functional networks related to cognitive, language, and motor development, ICA enables the model to better capture the neural patterns underlying developmental outcomes. This process not only preserves essential details but also reduces the computational complexity associated with high-dimensional raw data.
Furthermore, ICA’s ability to extract biologically meaningful components allows the model to focus on critical brain networks, improving accuracy and robustness over using the more variable and noisy raw data. This targeted approach highlights the value of combining neuroscience-driven feature engineering with advanced machine learning methods. When paired with multi-label learning, as shown in Table 6, ICA-extracted features yield the best results, demonstrating similar performance boosts to those observed with raw fMRI volumes. This underscores the effectiveness of integrating ICA into our pipeline to enhance predictive power and model efficiency.
Table 5: Mean performance with standard deviation through 5-cross-validation by using IC-extracted features, facilitating single-label predictions.
@@
Table 6: Mean performance with standard deviation through 5-cross-validation by using IC-extracted features, facilitating multi-label predictions.
@@ 3.5 Comparative Analysis
We have summarized our experimental results via boxplots in Figure 4, while omitting results obtained via transfer-learning and higher-dimensional ICs due to no statistically significant performance changes. Across all metrics, Multi-ICA demonstrates significant and consistent improvements over the best-performing baseline model for most tasks, particularly in Cognitive MAE (p = 0.004) and Motor MAE (p = 0.002). These results underscore the advantage of incorporating ICA-extracted features in the regression of cognitive and motor delays.
Similarly, for regression tasks, significant reductions in cognitive regression $\text{MSE}_{adj}$ (p = 0.036) and language regression $\text{MSE}_{adj}$ (p = 0.008) highlight the robustness of the Multi-ICA approach in modeling complex developmental outcomes. Importantly, these improvements hold across multiple target domains, reinforcing the utility of a multi-label learning paradigm.
In classification, Multi-ICA also achieves significant improvements in language classification accuracy (p = 0.004) and motor classification AUC (p = 0.044), while maintaining competitive performance across most other metrics. However, some metrics – such as cognitive classification AUC (p = 0.413) – did not show statistically significant differences compared to the baseline. These results suggest that the model’s performance in certain domains, particularly language-related tasks, may benefit from further domain-specific optimization.
Overall, Multi-ICA paired with multi-label learning consistently outperforms the strongest baseline model, confirming its utility for both regression and classification tasks in the context of predicting developmental outcomes.
@@ 4 Interpretation via Integrated Gradients
As in the original SwiFT study (Kim et al. 2023), we used Integrated Gradient with Smoothgrad sQuare (IG-SQ), implemented through the Captum framework (Rahman et al. 2022) (Adebayo et al. 2018) (Kokhlikyan et al. 2020). This method allowed us to identify spatial representations in brain regions that contribute to Bayley-III classification tasks. This analysis was based on the best-performing model, which combined multi-label prediction with ICA-extracted features. The use of independent components (ICs) facilitated neuroscientifically more useful interpretations by linking model outputs to known functional networks (Beckmann et al. 2009).
For each test sample, we generated 4D IG-SQ maps and excluded misclassified instances to ensure the validity of interpretations. The resulting attribution maps were normalized, Gaussian-smoothed, and averaged across subjects to identify spatial patterns consistently associated with cognitive, language, and motor outcomes. Additionally, we aggregated these maps across ICs to derive comprehensive interpretations of brain activity relevant to each developmental domain.
@@
The analysis revealed distinct brain regions whose spatiotemporal dynamics contribute to each developmental outcome (Figure 5). This data-driven approach reveals meaningful neuroscientific interpretations, demonstrating that regions within the distinct functional networks correspond closely to neurodevelopmental outcomes, reflecting their established roles in cognitive and neural development. For cognitive delay risk, activations were observed in the medial prefrontal cortex (coronal view), thalamocortical circuit (sagittal and axial view), and posterior parietal association cortex (sagittal view). These findings align with neuroscience literature: the identified brain regions are key to early sensory processing, attention, executive function, and spatial cognition, all foundational to normal cognitive development (Mair et al. 2022) (Benarroch and Benarroch 2021) (Shomstein 2012).
In Figure 6, network-level predictions are further explored ICA-extracted features. For instance, IC5 likely represents the Executive Control Network (ECN)(Li et al. 2019), a critical network for cognitive functions such as attention and decision-making. Its contribution to the overall interpretation, as shown in Figure 5, is evident through activations in the medial prefrontal cortex, a hallmark region of the ECN.
For linguistic delay risk (middle), activations in Wernicke’s area align with its well-documented role in language comprehension and development (Binder 2015). Similarly, for motor skill delays (bottom), activations in the primary motor cortex and supplementary motor areas (SMA) highlight their roles in gross and fine motor control and motor pattern learning, respectively (Graziano 2006; Tanji and Shima 1994).
@@
Figure 6: A glimpse into network-level interpretations, focusing on IC 5 (left), possibly representing the Executive Control Network (ECN). For predicting the risk of cognitive delay, can see how this brain-network adds to activation in the medial prefrontal cortex (right).
@@ 5 Concluding Remarks
In this study, we demonstrated that the SwiFT (Swin 4D fMRI Transformer) model provides a significant improvement in evaluating neonatal fMRI data for predicting neurodevelopmental outcomes. By integrating multi-label learning and leveraging ICA-extracted features, we achieved enhanced predictive accuracy while improving model interpretability. These advances suggest that SwiFT has the potential to play a pivotal role in the early detection of developmental delays, paving the way for personalized therapeutic interventions for at-risk neonates.
Despite these accomplishments, limitations remain. The imbalanced nature of the dataset poses a challenge, impacting the reliability of the classification tasks. While our approach partially mitigated this issue, future work should explore advanced strategies for handling data imbalance (Buda, Maki, and Mazurowski 2018). Additionally, further validation of ICA-extracted features as proxies for brain-network-level mechanisms is necessary to strengthen the biological interpretability of our findings. Also, expanding this framework to include datasets from other age groups, such as toddlers, children, and adults, could enhance the generalizability of the model. Finally, extending the multi-label learning paradigm to incorporate additional target variables, such as the Q-CHAT score for early autism screening, offers an exciting direction for future research.
In conclusion, this work establishes a robust foundation for advancing predictive and interpretable models of neurodevelopment. With continued refinement and access to diverse, large-scale datasets, the SwiFT Transformer holds significant potential for innovations in neuroscience and personalized medicine.

### v2

@@ 3.4.2 Results
For each dimensionality, the model was tuned using a predefined hyperparameter search space, including learning rate, patch size, and data augmentation strategies (see Section 3.3). With ICA reducing the input data to a fraction of its original dimensionality, careful tuning was critical to ensure effective convergence. Given the data demands of transformer architectures and the reduced training data size, we increased the maximum training epochs to 100. To mitigate overfitting, we applied data augmentation techniques such as affine and intensity transformations to enhance sample diversity and improve generalizability. Due to the specific sequence lengths required for IC-based training (42 and 100), transfer learning was not utilized in these experiments. We addressed potential data leakage by ensuring that the 100 healthy subjects used for Group ICA were always included in the training set, avoiding overlap with validation or test sets. This step minimized bias and maintained the reliability of our results. Experimental outcomes are summarized in Tables 3, presenting the average test performance and standard deviations of the fine-tuned models.
@@ Discussion
This study demonstrates that integrating Group ICA-based dimensionality reduction with SwiFT significantly enhances predictions of neurodevelopmental outcomes from neonatal fMRI data. By extracting biologically meaningful features via ICA and leveraging multi-label learning, the approach preserves critical neural information while reducing computational complexity. SwiFT effectively models local and global spatiotemporal patterns in 4D fMRI data, underscoring the synergy between neuroscience-driven feature engineering and advanced machine learning. As can be seen in Table 3, our results indicate that ICA dimensionality (25 vs. 100 ICs) does not substantially impact performance, suggesting that ICA robustly captures essential neural features regardless of the number of components. Furthermore, single-label learning with ICA-extracted features outperformed raw fMRI predictions, particularly for language classification and motor regression tasks, highlighting ICA’s ability to retain biologically meaningful information while reducing noise. The combination of ICA and multi-label learning further enhanced predictive power, aligning with previous findings on the interrelated nature of developmental outcomes. These findings reinforce ICA’s value as a preprocessing step, enabling the model to focus on key neural networks and improve accuracy and efficiency. This targeted approach demonstrates the potential of integrating neuroscience-driven features with attention-based architectures for advancing neurodevelopmental research and clinical applications.
Table 3: SwiFT’s performance with IC-extracted features as input on single-label (left) and multi-label (right) predictions.
@@ 3.5 Comparative Analysis
We have summarized our experimental results via boxplots in Figure 4, while omitting results obtained via transfer-learning and higher-dimensional ICs due to no statistically significant performance changes. After statistical significance analysis between the best-performing baseline model and SwiFT variation for each task across all metrics, Multi-ICA demonstrates significant and consistent improvements over the baseline model for most tasks, particularly in Cognitive MAE ($p=0.004$) and Motor MAE ($p=0.002$). Similarly, for regression tasks, significant reductions in cognitive regression $\text{MSE}_{adj}$ (p = 0.036) and language regression $\text{MSE}_{adj}$ (p = 0.008) highlight the robustness of the Multi-ICA approach in modeling complex developmental outcomes. Importantly, these improvements hold across multiple target domains, reinforcing the utility of a multi-label learning paradigm. In classification, Multi-ICA also achieves significant improvements in language classification accuracy (p = 0.004) and motor classification AUC (p = 0.044), while maintaining competitive performance across most other metrics. However, some metrics – such as cognitive classification AUC (p = 0.413) – did not show statistically significant differences compared to the baseline. These results suggest that the model’s performance in certain domains, particularly language-related tasks, may benefit from further domain-specific optimization. Overall, Multi-ICA paired with multi-label learning consistently outperforms the strongest baseline model, confirming its utility for both regression and classification tasks in the context of predicting developmental outcomes.
@@ 4 Interpretation via Integrated Gradients
As in the original SwiFT study [19], we used Integrated Gradient with Smoothgrad sQuare (IG-SQ), implemented through the Captum framework [28] [29] [30]. This method allowed us to identify spatial representations in brain regions that contribute to Bayley-III classification tasks. This analysis was based on the best-performing model, which combined multi-label prediction with ICA-extracted features. Leveraging ICs allowed for neuroscientifically meaningful interpretations by associating model outputs with known functional networks [40]. For each test sample, we generated 4D IG-SQ maps and excluded misclassified instances to ensure the validity of interpretations. The resultant maps were normalized, Gaussian-smoothed, and averaged across subjects to identify spatial patterns consistently associated with cognitive, language, and motor outcomes. Additionally, we aggregated these maps across ICs to derive comprehensive interpretations of brain activity relevant to each developmental domain.
Figure 5: A glimpse into network-level interpretations, focusing on IC 5 (left) possibly representing the Executive Control Network (ECN). We can see how this network adds to activation in the medial prefrontal cortex for prediction of cognitive delay (right).
The analysis revealed distinct brain regions whose spatiotemporal dynamics contribute to each developmental outcome (Figure 6). For cognitive delay risk, activations were observed in the medial prefrontal cortex (coronal view), thalamocortical circuit (sagittal and axial view), and posterior parietal association cortex (sagittal view). These findings align with neuroscience literature: the identified brain regions are key to early sensory processing, attention, executive function, and spatial cognition, all foundational to normal cognitive development [31] [32] [33]. In Figure 5, network-level predictions are further explored. For instance, IC5 likely represents the Executive Control Network (ECN)[34], a critical network for cognitive functions such as attention and decision-making. Its contribution to the overall interpretation, as shown in Figure 6, is evident through activations in the medial prefrontal cortex (mPFC), a hallmark region of the ECN. For the risk of linguistic delay (middle), activations in Wernicke’s area align with its well-documented role in language comprehension and development [35]. Similarly, for motor skill delays (bottom), activations in the primary motor cortex and supplementary motor areas (SMAs) highlight their roles in gross and fine motor control and motor pattern learning, respectively [36] [37].
@@ 5 Concluding Remarks
In this study, we demonstrated that SwiFT provides a significant improvement in evaluating neonatal fMRI data for predicting early neurodevelopmental outcomes. By integrating multi-label learning and leveraging ICA-extracted features, we achieved enhanced predictive accuracy while improving model interpretability. These advances suggest that SwiFT has the potential to play a pivotal role in the early detection of developmental delays, paving the way for personalized therapeutic interventions for at-risk neonates. The clinical relevance of such a model is strengthened by a study suggesting that therapeutic interventions to treat neurodevelopmental disorders may be more effective if done during the early stages of brain development [41]. Despite these accomplishments, limitations remain. The imbalanced nature of the dataset poses a challenge, impacting the reliability of the classification tasks. While our approach partially mitigated this issue, future work should explore advanced strategies for handling data imbalance, such as oversampling or even synthetic data generation. Additionally, further validation of ICA-extracted features as proxies for brain-network-level mechanisms is necessary to strengthen the biological interpretability of our findings. Also, expanding this framework to include datasets from other age groups, such as toddlers, children, and adults, could enhance the generalizability of the model. Extending the multi-label learning paradigm to incorporate additional target variables, such as the Q-CHAT score for early autism screening, offers an exciting direction for future research and could be implemented into the current pipeline without major changes Finally, pretraining on adult data could provide a robust foundation for the model, but has shown no generalizability to neonates in our experiments. Since the use of small neonatal datasets increases the risk of overfitting, examining other pretraining paradigms than contrastive learning, such as Masked Image Modeling, may be beneficial.
In conclusion, this work establishes a robust foundation for advancing predictive and interpretable models of neurodevelopment. With continued refinement and access to diverse, large-scale datasets, SwiFT holds significant potential for innovations in neuroscience and personalized medicine.